### v1

# Aligning MLLM Benchmark With Human Preferences via Structural Equation Modeling

Abstract
Evaluating multimodal large language models (MLLMs) remains a fundamental challenge due to a lack of structured, interpretable, and theoretically grounded benchmark designs. Existing benchmarks often adopt heuristic-based task groupings with unclear cognitive targets, thus resulting in overlapping abilities, redundant indicators, and limited diagnostic power. In this work, we propose a novel framework for aligning MLLM benchmark based on Structural Equation Modeling (SEM) to analyze and quantify the internal validity, dimensional separability, and contribution of benchmark components. Motivated by the observed limitations of current designs, we further introduce a novel capability hierarchy grounded in Piaget’s theory of cognitive development, dividing MLLM abilities into three hierarchical layers, i.e., Perception, Memory, and Reasoning. We reorganize existing MLLM benchmarks under the proposed framework and construct a new benchmark named Gold. Experimental results demonstrate that the proposed benchmark exhibits stronger interpretability, reduced indicator redundancy, and clearer cognitive consistency compared to existing approaches. Our benchmark is available at https://github.com/tianyu-zou/GoldBench.
1 Introduction
The rapid advancements in the field of multimodal learning have been driven by the emergence of increasingly Multimodal Large Language Models (MLLMs). From OpenAI’s GPT-4o achiam2023gpt and Google’s Gemini team2023gemini to a variety of open-source models such as LLaVA liu2023visual ; touvron2023llama , mPLUG-Owl ye2023mplug , and BLIP-2 li2023blip , these models exhibit remarkable capabilities in integrating and processing multiple modalities, including text, images, and videos. As MLLMs continue to evolve, the need for comprehensive evaluation frameworks becomes increasingly critical to assess their reasoning abilities, multimodal understanding, and generalization performance fu2024mme ; li2024embodied . A rigorous evaluation is essential not only for benchmarking advancements in model development but also for identifying potential biases, addressing ethical concerns, and ensuring the responsible and reliable deployment of these models across various real-world applications feng2025benchmarking ; qin2025survey ; zhang2024q .
Evaluations not only reveal the advantages and shortcomings of these models across various tasks and dimensions, but they also provide clear directions for future enhancements zhu2023large ; li2024survey ; huang2024survey . Currently, the evaluation of MLLMs primarily relies on benchmark datasets composed of large-scale question-answer pairs guo2025r ; li2025benchmark . According to statistics from Opencompass 2023opencompass , there are now over 300 evaluation benchmarks for MLLMs, with more than 100 of them integrated into various multimodal evaluation suites. However, the rapid proliferation of evaluation benchmarks has also introduced new challenges: the lack of unified standards makes it difficult to assess the quality, coverage, and reliability of these benchmarks li2025information ; zhang2025redundancy . As a result, it becomes increasingly hard to determine which benchmarks offer meaningful insights into model capabilities and which may lead to biased or misleading conclusions.
A particularly notable trend in recent benchmarks is the widespread adoption of tiered capability frameworks, that is attempts to deconstruct model performance into structured dimensions. We have found that while most evaluation frameworks embrace this approach, there exist significant differences among them in terms of capability dimensions and task design. Typical evaluation benchmarks include: SEED-Bench li2024seed , which proposes a five-level capability framework and constructs 27 tasks; MV-Bench li2024mvbench , which divides 47 evaluation tasks based on six core capabilities; and MMT-Bench ying2024mmt , which expands capability dimensions to 32 with an accompanying 162 fine-grained tasks. This phenomenon of ”dimension expansion” reflects the ongoing academic exploration of model understanding. However, it also raises a fundamental methodological concern: How can we construct a theoretically grounded and empirically verifiable capability structure that avoids arbitrary task proliferation and ensures interpretability?
To address this, we elaborate a novel framework by using Structural Equation Modeling (SEM) ullman2012structural as a principled approach to define and validate the latent structure of model abilities based on observed task performance. Recognizing human evaluation as the gold standard for reliability, its extensive labor cost and limited scalability create an urgent need for an alternative that approximates human judgment while dramatically reducing manual effort. The proposed framework meets this requirement by modeling the causal relationships between observed task indicators and the unobservable constructs they form. It organizes characteristics such as perception and cognition into empirically testable measurement models. Unlike traditional statistical or heuristic analyses, it not only infers latent variables from observable data but also reveals the causal contributions of individual features to overall evaluation quality, and can be calibrated to align with human cognitive preferences. By adopting the proposed framework for the evaluation of MLLM benchmarks, we can systematically assess the structural validity and practical effectiveness of the benchmark.
To further leverage our evaluation framework, we draw on Piaget’s theory of cognitive development piaget2000piaget , using its account of how humans acquire perceptual, mnemonic, and reasoning abilities to derive the human-centered Gold benchmark. Traditional benchmarks, driven largely by heuristic task groupings, often yield inconsistent or overlapping ability hierarchies that neither reflect human cognition nor guide task design in a principled way. A Piaget-inspired bottom-up stratification offers a clear remedy. By aligning MLLM capabilities with the Sensorimotor, Preoperational, and (Concrete & Formal) Operational stages of human development, we define three continuous ability layers in Gold: Perception, Memory, and Reasoning. This human-cognitive foundation enhances interpretability and consistency in benchmarks.
Figure 1: An architecture of SEM-based evaluation framework.
Our main contributions are as follows:
  • SEM-based Evaluation Framework: To the best of our knowledge, we for the first time apply SEM in MLLM assessment, which models the causal relationships between observable task metrics and latent capabilities. This approach not only introduces a novel evaluation pipeline but also ensures that it provides a transparent, verifiable, and reproducible framework for MLLM benchmark evaluation.
  • Piaget-Inspired Hierarchical Capability Structure: By leveraging Piaget’s theory of cognitive development, we define a three-tier capability hierarchy, i.e., Perception, Memory, and Reasoning, which helps us systematically organize tasks, mitigate dimension inflation, and enhance interpretability.
  • Gold Benchmark: We propose the human-centered Gold benchmark, derived from our Piaget-inspired hierarchical framework, to systematically evaluate MLLM across perceptual, mnemonic, and reasoning tasks in a structured, cognitively grounded manner.
Figure 2: Structural analysis of MMBench by using SEM-based framework. MMBench shows stronger task contributions but suffers from redundancy.
2 Evaluating Benchmark Structures via SEM
2.1 Revisiting of Structural Equation Modeling
SEM can be constructed based on causal measurement models, organizing multiple observed variables into a coherent structural framework. It consists of two core components: the measurement model, which defines the relationships between observed variables and their underlying latent constructs, and the structural model, which captures the causal relationships among latent variables. As discussed in Appendix, we adopt a formative measurement model as shown in Figure 1. The measurement model specifies the relationships between observed and latent variables as follows: $X=\Lambda_{x}\xi+\delta$, $Y=\Lambda_{y}\eta+\epsilon$, where $X$ and $Y$ are observed exogenous and endogenous variables, $\xi$ and $\eta$ are latent variables, $\Lambda_{x}$ and $\Lambda_{y}$ are factor loading matrices, and $\delta$, $\epsilon$ are measurement errors. The structural model describes causal relations among latent variables: $\eta=B\eta+\Gamma\xi+\zeta$, with $B$ and $\Gamma$ representing path coefficients and $\zeta$ the residuals. Together, the measurement and structural models provide a comprehensive framework for assessing both direct and indirect effects.
2.2 Structural Analysis of Existing Benchmarks
Interestingly, we further apply the proposed SEM-based evaluating framework to analyze existing MLLM benchmarks. Specifically, we investigate whether the task configurations within current benchmarks provide complementary information while jointly forming meaningful latent constructs. Our analysis reveals that most benchmarks struggle to balance low redundancy with high construct validity.
To illustrate this issue in detail, we conduct a case study on two representative benchmarks: MMBench liu2024mmbench and BLINK fu2024blink . Leveraging the performance reports of over 190 MLLMs collected via the VLMEvalKit duan2024vlmevalkit toolkit, we construct formative structural equation models based on the hierarchical task taxonomy specified by each benchmark. In our modeling framework, individual tasks serve as observed indicators, which are grouped into intermediate latent variables representing ability dimensions such as perception, reasoning, and cross-modal understanding. These first-order latent abilities are then aggregated into a second-order latent variable that captures the overall capability of an MLLM. Crucially, to align this latent construct with human-perceived performance, we incorporate the Chatbot Arena chiang2024chatbot score as an external observed indicator of the top-level latent variable. This hierarchical SEM structure enables us to evaluate how well benchmark-defined ability dimensions, as derived from task-level scores, explain human-judged overall performance.
In the case of MMBench, as shown in Figure 2(a), tasks exhibit high contribution to their corresponding ability dimensions, with outer loadings exceeding 0.85 for most tasks. However, the analysis of Variance Inflation Factor (VIF) values in Figure 2(b) shows a concerning level of multicollinearity, particularly for tasks like Image Scene (VIF = 13.96) and Attribute Recognition (VIF = 6.12), indicating a high degree of redundancy within the benchmark. This redundancy undermines the overall efficiency of the measurement model, as the tasks are provide insufficiently independent information.
On the other hand, BLINK as shown in Figure 3 demonstrates a different issue. While the VIF values for BLINK are generally low (with a maximum value of 3.09 for Art Style), indicating that task redundancy is less of a concern, the outer loadings of several tasks are lower compared to MMBench. Tasks such as IQ Test (outer loading = 0.354) and Counting (outer loading = 0.686) contribute less to their respective constructs, suggesting that the benchmark lacks sufficiently strong task performance indicators for each cognitive ability dimension.
Figure 3: Structural analysis of BLINK by using SEM-based framework. BLINK has lower redundancy but weaker task contributions.
Despite structural differences, both MMBench and BLINK exhibit limited discriminant validity among their ability dimensions, as indicated by their respective HTMT matrices. Several HTMT values exceed or approach the 0.90 threshold, particularly between conceptually adjacent dimensions, revealing substantial overlap and suggesting that current ability taxonomies in both benchmarks require further conceptual clarification.
Overall, while MMBench and BLINK differ in their structural characteristics, MMBench showing strong but redundant task contributions, and BLINK presenting low redundancy but weaker indicators. Both suffer from poor discriminant validity in their ability dimensions. These findings expose fundamental limitations in current benchmark design and underscore the need for a more systematic approach to evaluating benchmark structure. These findings reveal the urgent need for a principled framework to evaluate the internal structure of MLLM benchmarks and to redefine ability dimensions based on sound cognitive foundations.
2.3 A SEM-based Evaluation Framework for Benchmark
To evaluate the robustness and interpretability of MLLM Benchmarks, we propose an evaluation framework for benchmark, which conducts a comprehensive concept-level analysis of the latent ability dimensions and their associated task indicators. It focuses on three key aspects: dimensional diversity, task contribution and indicator validity. These factors reflect the conceptual soundness and representational adequacy of the evaluation benchmark, as well as the redundancy and contribution of its indicators to the corresponding tasks.
Dimensional Diversity.
Dimensional diversity aims to quantify the discriminant validity among latent ability dimensions, thereby characterizing their conceptual independence and potential redundancy. Although existing benchmarks effectively capture multiple facets of MLLMs’ capabilities, overlaps between different dimensions may still exist. To rigorously assess discriminant validity, we adopt the Heterotrait–Monotrait Ratio of correlations (HTMT). For any two latent dimensions $\xi_{i}$ and $\xi_{j}$, the HTMT index is defined as
$\text{HTMT}_{ij}=\frac{\frac{1}{|H_{ij}|}\sum\limits_{(x_{p}\in H_{i},x_{q}\in H_{j})}\text{cor}(x_{p},x_{q})}{\frac{1}{|M_{ii}|}\sum\limits_{(x_{m},x_{n}\in H_{i})}\text{cor}(x_{m},x_{n})},$ (1)
where $H_{i}$ and $H_{j}$ denote the sets of observed indicators corresponding to latent dimensions $\xi_{i}$ and $\xi_{j}$, respectively. The set $H_{ij}$ comprises all possible pairs $(x_{p},x_{q})$ such that $x_{p}\in H_{i}$ and $x_{q}\in H_{j}$, representing heterotrait–heteromethod comparisons. Each $x_{p}$ and $x_{q}$ denotes an observed variable (e.g., a task-specific score) that loads onto the latent dimensions $\xi_{i}$ and $\xi_{j}$, respectively. Conversely, the set $M_{ii}$ includes all within-construct pairs $(x_{m},x_{n})$ such that both $x_{m}$ and $x_{n}$ belong to $H_{i}$, representing monotrait–heteromethod comparisons. The function $\mathrm{cor}(\cdot,\cdot)$ denotes the Pearson correlation coefficient between observed variables.
Empirical guidelines suggest that an HTMT value exceeding 0.90 indicates deficient discriminant validity and substantial conceptual redundancy between constructs, whereas lower values imply satisfactory separation. Building on this heuristic, we define the dimensional diversity score as
$D_{\mathrm{div}}=\frac{1}{2*\max_{i\neq j}\mathrm{HTMT}_{ij}}$ (2)
where $\max{i\neq j}\mathrm{HTMT}{ij}$ represents the maximum HTMT value across all distinct pairs of latent dimensions. The denominator is scaled by 2 to ensure that $D{\mathrm{div}}\in[0,1]$, given that HTMT values are typically bounded within the empirical range of [0, 1]. A higher $D_{\mathrm{div}}$ indicates stronger discriminant validity and thus greater conceptual distinctiveness among latent constructs.
Task Contribution.
Task contribution quantifies the relative importance of each task indicator in shaping its corresponding latent construct. In order to capture this notion in a formative measurement context, we adopt the concept of out-loading, which represents the standardized weight of each observed task on the latent dimension. To evaluate the contribution of tasks at the level of the entire benchmark, which consists of multiple ability dimensions $\xi_{=}\{\xi_{1},\dots,\xi_{m}\}$, we define the benchmark-level task contribution score as
$\begin{split}TC&=\frac{1}{\sum_{i=1}^{n}m_{i}}\sum_{i=1}^{m_{i}}\sum_{j=1}^{n}\left|\lambda_{i,j}\right|,\end{split}$ (3)
the metric $TC\in[0,1]$ provides a global estimate of task-level informativeness across the entire benchmark, with higher values indicating stronger and more concentrated contributions from individual tasks.
Indicator Validity.
Indicator validity is assessed both substantively and statistically, with the latter focusing on multicollinearity diagnostics among task indicators. For each indicator $x_{i,j}$ associated with latent dimension $\xi_{j}$, we compute the Variance Inflation Factor (VIF) by
$\mathrm{VIF}_{x_{i,j}}=\frac{1}{1-R^{2}_{x_{i,j}}},$ (4)
where $R^{2}_{x_{i,j}}$ is the coefficient of determination obtained by regressing $x_{i,j}$ on the remaining tasks. A higher VIF value indicates stronger collinearity, suggesting that the task may not provide unique information beyond what is already captured by others.
To consolidate individual VIF values into a single benchmark-level measure, we define the composite validity score $D_{\mathrm{valid}}$ as the inverse of the geometric mean of all task-level VIFs
$D_{\mathrm{valid}}=\left(\prod_{j=1}^{n}\mathrm{VIF}_{x_{j}}\right)^{-1/n},$ (5)
where $n$ denotes the total number of task indicators in the benchmark. This aggregate score $D_{\mathrm{valid}}\in(0,1]$ reflects the overall degree of indicator independence across the benchmark. Lower values indicate potential redundancy and multicollinearity issues at a global level, while higher values suggest stronger statistical validity of the task indicators as formative constructs.
3 Gold Benchmark
3.1 Hierarchical Structure Based on Piaget’s Developmental Theory
We propose a new capability hierarchy grounded in Jean Piaget’s theory of cognitive development, a well-established framework in developmental psychology. Unlike previous heuristic-based groupings, the proposed approach introduces a human-centered, theory-driven structure to model the cognitive underpinnings of MLLM performance. By anchoring the evaluation framework in developmental principles, we aim to achieve a more interpretable, hierarchical, and cognitively faithful representation of model abilities.
This Piaget-inspired cognitive structure is embedded into our benchmark design, with each task aligned to one of the three hierarchical cognitive layers (Perception, Memory, and Reasoning). SEM confirms the distinctiveness and validity of this stratification, demonstrating improved dimensional separability, reduced task redundancy, and stronger measurement consistency compared to conventional heuristic groupings.
3.2 Gold Construction
Algorithm 1 Cognitive Benchmark Construction via Piaget Theory and SEM
Hyperparameters: VIF threshold $\delta_{\text{VIF}}$, loading threshold $\lambda_{\min}$
Input: Normalized Task indicators $\mathcal{X}$ from various benchmarks
Step 1: Define latent variables: Perception, Memory, Reasoning
Step 2: Assign tasks $x_{i}\in\mathcal{X}$ to latent constructs
Step 3: Initialize weights $w\leftarrow 1$
while there exists task with VIF $>\delta_{\text{VIF}}$ or outer loading $<\lambda_{\min}$ do
     Estimate model via PLS-SEM (maximize explained variance)
     Remove tasks with VIF $>\delta_{\text{VIF}}$ (redundancy)
     Remove tasks with outer loading $<\lambda_{\min}$ (low quality)
end while
Step 3: Validate model with Cronbach’s $\alpha$ and HTMT
Output: A theory-aligned benchmark structure with validated task indicators
To construct a human-aligned benchmark structure, we map tasks from existing multimodal datasets to cognitive processes defined in a Piaget-inspired framework (Section 3.1). This process is formalized by using SEM, as outlined in Algorithm 1. We begin by curating an indicator pool by collecting tasks from four widely adopted multimodal benchmarks: MME fu2023mme , MMMU-Bench yue2024mmmu , BLINK fu2024blink and MMBench liu2024mmbench . Each task is manually annotated and assigned to one cognitive layer based on its processing demands (e.g., OCR $\rightarrow$ Perception; landmark recall $\rightarrow$ Memory; domain-specific inference $\rightarrow$ Reasoning).
Each task is manually annotated and aligned to one of the three cognitive domains based on its processing demands. We then apply Partial Least Squares SEM (PLS-SEM) to model the relationships among latent constructs and their associated indicators. This method is well-suited to our context, where sample size is limited and indicators may be collinear.
To refine the model, we iteratively prune the indicator pool by removing tasks exhibiting high multicollinearity (variance inflation factor $>\delta_{\text{VIF}}$) or insufficient indicator quality (outer loading $<\lambda_{\min}$). Following prior empirical standards fong2013hair ; 10990127 , we set $\delta_{\text{VIF}}=5$ and $\lambda_{\min}=0.75$. This process enhances both statistical validity and conceptual coherence. We further evaluate construct reliability and discriminant validity using Cronbach’s $\alpha$ and the heterotrait–monotrait (HTMT) ratio.
The final output, Gold, is a compact and theory-grounded task set with optimized indicator weights and path coefficients. Each selected task robustly represents its assigned cognitive function, enabling interpretable, scalable, and semantically aligned evaluation of MLLMs.
3.3 Statistics.
Figure 4: Structural analysis of Gold by using SEM-based framework.
The final Gold Benchmark contains 12 tasks, selected from an initial pool of 85 across the four benchmarks. Our analysis follows the guidelines proposed by J. Hair hair2019use . For measurement model, we examine the indicator weights to evaluate their contribution to the latent constructs. As depicted in the figure, the weights are consistently greater than 0.75. This outcome indicates that the retained indicators possess strong explanatory power and are effectively capturing the underlying constructs they are intended to measure.
Table 1: Reliability and Validity Metrics for Latent Constructs
Construct Cronbach’s $\alpha$ Reliability Composite reliability Convergent validity
Memory 0.702 0.732 0.823 0.607
Perception 0.804 0.813 0.885 0.720
Reasoning 0.902 0.906 0.927 0.719
To further verify the statistical quality of the benchmark, we conduct standard reliability and validity tests on the latent constructs. As shown in Table 1, all constructs meet the recommended thresholds for internal consistency and convergent validity. Specifically, Cronbach’s $\alpha$ values are all above 0.70, reflecting acceptable internal reliability. Composite reliability scores exceed 0.80 across all dimensions, confirming the coherence of the grouped indicators. In terms of convergent validity, all constructs exhibit Average Variance Extracted (AVE) values above 0.60, suggesting that each construct explains a substantial portion of the variance in its associated indicators.
We also assess multicollinearity using the Variance Inflation Factor (VIF). As shown in Figure 4(b), all indicators yield VIF values well below the conservative threshold of 5 (ranging from 1.12 to 2.75), suggesting the absence of problematic multicollinearity and supporting the discriminant structure of the model. Finally, we evaluate discriminant validity using the Heterotrait-Monotrait (HTMT) ratio in Figure 4(c). All pairwise HTMT values are below the recommended cutoff of 0.90, with the highest observed value being 0.86 (between Memory and Perception). This result confirms that each latent construct captures a distinct cognitive domain and is not confounded with others.
These statistical results collectively validate the robustness of our benchmark construction. They demonstrate that the selected tasks reliably capture the latent cognitive dimensions defined in our Piaget-inspired framework, thus ensuring both empirical rigor and theoretical alignment.
4 Experimental Study
4.1 Experiment Setting
To evaluate the quality and interpretability of the Gold benchmark, we design a series of experiments involving both open-source and closed-source MLLMs. Our evaluation includes two core settings: (1) Main Experiment, where we assess mainstream models using a standardized pipeline to reveal overall performance differences; (2) SEM-based Benchmark Evaluation Experiment, where we validate Gold’s diagnostic effectiveness by measuring key metrics such as modality diversity ($D_{div}$), task contribution ($TC$), and diagnostic validity ($D_{valid}$). We further assess its interpretability by comparing these outputs with human ratings to examine alignment with human judgment.
Evaluation Models. For main experiment, we evaluate both closed-source and open-source MLLMs. For the closed-source group, we include GPT-4o achiam2023gpt . For the open-source group, we consider a range of models with varying scales, including Qwen2.5-VL bai2025qwen2 , Gemma3 team2025gemma , and InternVL3 zhu2025internvl3 . These models represent the current frontier of multimodal LLM development and cover diverse architecture families and parameter sizes. For the benchmark evaluation experiment, we leverage evaluation results from 165 multimodal language models collected via VLMEvalKit duan2024vlmevalkit , which provides standardized testing across multiple benchmarks. The Metrics used in the experiment will be described in detail in the Appendixes.
Table 2: Comparison of state-of-the-art methods on the Gold. BMK and CS denote Basic Medical Knowledge and Computer Science, respectively. Overall scores are computed by weighting capability-level scores according to their corresponding task contributions.
Methods Model size Overall Perception Memory Reasoning
Color Count OCR Artwork Landmark BMK Biology CS Economics Electronics Math
MLLM (closed source)
[Uncaptioned image]  GPT-4o - 165.7 92.5 92.5 96.3 72.6 91.0 73.3 56.7 73.3 80.0 76.7 70.0
Open source
[Uncaptioned image]  Gemma3 4B 118.6 54.2 55.0 81.3 62.5 80.1 53.3 36.7 46.7 46.7 53.3 50.0
[Uncaptioned image]  Gemma3 12B 139.4 75.0 75.0 85.0 75.6 88.3 63.3 33.3 60.0 73.3 63.3 53.3
[Uncaptioned image]  InternVL3 2B 116.9 86.7 77.5 85.0 78.1 84.5 30.0 40.0 40.0 56.7 50.0 40.0
[Uncaptioned image]  InternVL3 9B 153.6 89.2 86.7 96.3 80.9 87.9 63.3 66.7 63.3 60 73.3 56.7
[Uncaptioned image]  InternVL3 38B 163.7 86.7 82.5 96.3 82.4 92.3 76.7 50.0 70.0 76.7 83.3 63.3
[Uncaptioned image]  InternVL3 78B 172.9 91.7 85.0 96.3 82.1 92.9 83.3 66.7 80.0 80.0 76.7 66.7
[Uncaptioned image]  Qwen2.5-VL 3B 113.7 85.0 85.0 66.3 72.3 90.1 66.7 43.3 50.0 40.0 53.3 26.7
[Uncaptioned image]  Qwen2.5-VL 7B 148.2 92.5 82.5 96.3 147.5 73.8 91.6 43.3 56.7 63.3 76.7 63.3
[Uncaptioned image]  Qwen2.5-VL 72B 157.6 95.0 86.65 96.3 70.4 79.8 73.3 60.0 63.3 80.0 83.3 50.0
4.2 Main Results
In this section, we report the evaluation outcomes of various open-source and closed-source MLLMs on the Gold benchmark under the Main Experiment setting. Table 2 presents the detailed results, where models are assessed across three cognitive layers—Perception, Memory, and Reasoning—further subdivided into interpretable subtasks such as color recognition, OCR, basic medical knowledge, and domain-specific logical reasoning.
Overall Performance. GPT-4o, the only closed-source model in our comparison, achieves strong performance with an overall score of 165.7, particularly excelling in perception and memory tasks. However, top-tier open-source models such as InternVL3-78B (172.9) and Qwen2.5-VL-72B (157.6) actually surpass or closely approach GPT-4o in overall capability, suggesting that the performance gap between proprietary and open-source MLLMs is narrowing.
Capability-Level Comparison. Across all models, Perception tasks (Color, Count, OCR) exhibit the highest average scores, indicating that current MLLMs are well-optimized for low-level visual understanding. In contrast, Reasoning tasks—especially those requiring domain-specific knowledge (e.g., Electronics, Math, Economics) are the most challenging, with wide performance variability. This suggests that abstract and high-order reasoning remains a major bottleneck for multimodal intelligence.
Model Size and Performance. Larger model variants generally perform better, especially within the InternVL3 and Qwen2.5-VL families. For example, InternVL3 improves from 116.9 (2B) to 172.9 (78B), showing consistent gains across all layers. However, despite these improvements, performance on certain reasoning subfields (e.g., electronics, economics) still lags, indicating that increased model size alone does not fully bridge capability gaps.
4.3 SEM-based Benchmark Evaluation Result Analysis
As shown in Table LABEL:tab:diagnostic_results, Gold achieves the best balance among the three evaluation dimensions, with the highest average score (0.656) across all metrics. It combines broad cognitive coverage, strong task informativeness, and robust statistical validity, demonstrating the strength of theory-driven benchmark restructuring under the SEM-based framework. In contrast, some existing benchmarks show localized strengths but lack holistic consistency. Moreover, Gold shows the highest alignment with human ratings (Pearson $r=0.7359$), reflecting its superior interpretability and evaluation reliability.
Table 3: Diagnostic metrics for existing MLLM benchmarks across four dimensions.
Benchmark $D_{div}$ $TC$ $D_{valid}$ Avg
SEEDBenchV2 0.605 0.871 0.081 0.519
MVBench 0.094 0.584 0.522 0.400
MMEBench 0.549 0.861 0.420 0.610
MMMUBench 0.492 0.808 0.442 0.581
BLINK 0.414 0.715 0.618 0.582
QBench 0.513 0.926 0.246 0.562
MMBench 0.465 0.920 0.252 0.546
Gold (Ours) 0.579 0.833 0.557 0.656
Figure 5: Correlation with human evaluation across benchmarks.
Dimensional Diversity. Dimensional diversity ($D_{\text{div}}$) measures how distinct and non-overlapping the latent ability dimensions are. It is computed as the inverse of the highest Heterotrait–Monotrait (HTMT) ratio across all construct pairs. Benchmarks like SEEDBenchV2 (0.605), MMEBench (0.549), and our Gold (0.579) exhibit strong discriminant validity across dimensions, reflecting diverse and well-separated cognitive constructs. In contrast, MVBench (0.094) shows the lowest $D_{\text{div}}$, suggesting highly entangled constructs that may compromise diagnostic clarity.
Task Contribution. Task contribution ($TC$) assesses the average magnitude of standardized loadings for all indicators across their respective constructs. High $TC$ values indicate that tasks meaningfully shape latent abilities with low redundancy. Most benchmarks perform well on this metric, including QBench (0.926), MMBench (0.920), and MMEBench (0.861). Our Gold benchmark (0.833) maintains a high $TC$ while avoiding excessive task overlap. In contrast, MVBench again underperforms (0.584), indicating weaker task-level informativeness and possible construct inflation due to redundant indicators.
Indicator Validity. Indicator validity ($D_{\text{valid}}$) evaluates how independent task indicators are in terms of multicollinearity. Computed as the inverse geometric mean of their VIFs, higher scores imply stronger statistical uniqueness across tasks. BLINK (0.618), Gold (0.557), and MVBench (0.522) exhibit the highest indicator validity, suggesting that their tasks are statistically independent and less redundant. Meanwhile, SEEDBenchV2 (0.081) and QBench (0.246) show weak indicator validity, indicating significant multicollinearity among tasks that may hinder interpretability.
Correlation Analysis We evaluated the Pearson correlation between benchmark scores and human evaluation results. Among all compared benchmarks, Gold (ours) achieves the highest correlation (0.7359), demonstrating its superior alignment with human judgment. MMMU also shows strong consistency (0.702), whereas BLINK and MVBench exhibit weaker correlations (0.2746 and 0.369), suggesting that certain benchmarks may not effectively reflect real-world reasoning capabilities.
5 Conclusion
We propose a statistics-theory-driven evaluation framework for MLLMs, grounded in Piaget’s cognitive development theory. By structuring model capabilities into perception, memory, and reasoning, we construct a new benchmark, Gold, to enable fine-grained, interpretable, and statistically valid evaluation. Experimental results show that Gold achieves better indicator validity, lower redundancy, and higher cognitive alignment compared to existing benchmarks, demonstrating the effectiveness of our framework.
Appendix A Technical Appendices and Supplementary Material
A.1 Related Work
A.1.1 MLLM Benchmarks
A variety of benchmark datasets have been proposed to evaluate the capabilities of multimodal large language models (MLLMs), covering visual perception, language understanding, external knowledge integration, and complex reasoning. Early datasets such as COCO [37], VQA [38], and Visual7W [39] mainly assess perception and basic question answering abilities. Subsequent benchmarks like GQA [40] and VCR [41] emphasize compositional and commonsense reasoning, while OKVQA [42] and TextVQA [43] introduce challenges involving external knowledge and OCR-based understanding. ScienceQA [44] further targets scientific reasoning in educational contexts. As MLLMs develop, these earlier benchmarks are no longer sufficient to capture their increasingly diverse capabilities. Recent efforts, such as MMBench [25], OmniBench [45], and UniBench [46], aim to provide broader evaluations that include mathematics, medical diagnosis, chart understanding, and instruction following. However, these benchmarks typically rely on heuristic or empirical task categorization, leading to overlapping definitions, inconsistent classification, and limited diagnostic granularity, thus hindering efficiency analysis of model capabilities.
A.1.2 Evaluation for Benchmarks
In addition to evaluate the capabilities of MLLMs themselves, recent studies have begun to examine the design quality and diagnostic effectiveness of evaluation benchmarks. Most existing approaches follow a model-centric paradigm, where the quality of a benchmark is assessed indirectly through model performance across various tasks. Representative methods such as LIME [47] and CVR [48] incorporate human annotations to interpret model errors and seek a balance between task difficulty and error susceptibility. However, these ‘Human Eval’ approaches suffer from high labor costs and limited scalability. To address this limitation recent studies like Zerobench [49] and EnigmaEval [50] have explored ‘model Eval’ evaluation, which infers task difficulty and redundancy directly from model behavior. While more scalable, these methods risk introducing model-specific biases that compromise objectivity. Concurrently, efforts such as Information Density [18] investigate properties like fallacy, difficulty, redundancy, and diversity from a data-centric perspective, and Redundancy Principles [19] offering new insights for benchmark design. Nonetheless, these approaches often rely on heuristic analyses and lack principled theoretical grounding.
To bridge this gap, we propose a SEM-based framework that captures latent ability structures through causal modeling, enabling systematic evaluation of a benchmark’s construct validity and cognitive alignment. Building on this, we further introduce the Gold inspired by Piaget’s theory of cognitive development [24], organizing MLLM capabilities into a three-tier structure: Perception, Memory, and Reasoning, to enhance interpretability and theoretical coherence in benchmark design.
A.2 SEM-based Evaluation Framework Supported Paradigms
SEM can be categorized into formative models and reflective models based on the relationship between observed indicators and latent variables. In a reflective model, the latent variable is considered the source of variation in the observed variables, which are reflections of the latent variable. In contrast, a formative model assumes that the latent variable is "formed" by multiple observed indicators, and its value depends on the composition of these indicators.
In the evaluation framework for MLLMs, we focus on the model’s performance across various specific tasks, and these task scores collectively determine our judgment of a particular latent ability. In other words, a certain ability does not preexist and dominate the performance of individual tasks, but is instead constituted by the aggregation of task performance. This logic of "multiple observed indicators forming a latent variable" aligns clearly with the definition of a formative measurement model.
Therefore, in constructing the SEM for evaluating MLLM capabilities, it is more appropriate to employ a formative measurement model. Unlike reflective models that assume latent variables give rise to observed indicators, the formative approach better captures the nature of MLLM evaluation, wherein diverse task-specific scores contribute to the formation of underlying cognitive abilities. As shown in Figure 1, the formative measurement model specifies that each latent ability dimension $\xi_{j}$ is formed as a linear combination of its automatic task-level indicators, i.e., $\xi_{j}=\sum_{m_{j}}\lambda_{x_{m_{j},j}}x_{m_{j},j}$, where $x_{m_{j},j}$ denotes the performance score of the MLLM on the $m_{j}$-th task corresponding to the $j$-th ability, and $\lambda_{x_{m_{j},j}}$ represents the factor loading that quantifies the contribution of each task to the latent construct. Furthermore, the overall capability of the MLLM, denoted by $Y$, is conceptualized as a second-order latent construct synthesized from the individual ability dimensions, expressed as $Y=\sum_{i}\omega_{i}\xi_{i}$, where $\omega_{i}$ denotes the weight assigned to the $i$-th ability dimension. This hierarchical formulation enables a comprehensive and theoretically grounded assessment of MLLM performance, accounting for both granular ability-level indicators and the integrated overall capability.
A.3 More details about Hierarchical Structure Based on Piaget’s Developmental Theory
Concretely, inspired by the major stages of Piaget’s theory, we define three core layers of MLLM capabilities:
  • Perception Layer (Sensorimotor Stage): This layer captures low-level perceptual and sensory processing abilities, including visual parsing, object recognition, attribute identification, and vision-language spatial alignment. It tests whether the model can robustly process and ground raw multimodal inputs.
  • Memory Layer (Preoperational Stage): This layer focuses on symbolic representation and contextual memory, encompassing tasks such as long-range dependency tracking, metaphor comprehension, and situational recall. It assesses the model’s ability to retain, represent, and retrieve information in dynamic or abstract contexts.
  • Reasoning Layer (Concrete and Formal Operational Stages): This layer evaluates high-level cognitive capabilities, such as causal inference, multi-hop reasoning, mathematical problem-solving, and abstract cross-modal generalization. Tasks in this category are designed to test whether the model can perform complex, multi-step reasoning and apply knowledge flexibly to novel problems.
A.4 Limitations
While our evaluation framework offers a theory-driven and cognitively structured approach to MLLM assessment, it is not without limitations. The current version of the Gold benchmark, though aligned with Piaget’s cognitive development theory and effective in capturing core capabilities across perception, memory, and reasoning, may be overly simplistic for the latest generation of MLLMs. Many state-of-the-art models have demonstrated near-perfect performance on several sub-tasks, indicating a potential ceiling effect that limits the benchmark’s discriminative power. This also reflects the broader challenge of constructing cognitively grounded yet sufficiently challenging tasks that can meaningfully differentiate highly capable models. As part of future work, we plan to iteratively enrich the benchmark with more complex, ambiguous, and cognitively demanding items to better reflect real-world task diversity and the advancing frontiers of MLLM capabilities.
A.5 Metrics
Overall Metric.
To evaluate overall model performance, we adopt a weighted aggregation scheme in which each subtask score $s_{i}$ is multiplied by its corresponding task contribution weight $w_{i}$. These weights are derived from our structural equation modeling (SEM)-based framework, which captures the causal relationships between observed task performance and latent cognitive capabilities. This ensures that the final score reflects not only raw accuracy but also the theoretical importance of each cognitive dimension. The overall score is computed as:
Overall $\text{Overall Score}=\frac{\sum_{i=1}^{N}w_{i}\cdot s_{i}}{\sum_{i=1}^{N}w_{i}}.$ (6)
To support detailed capability analysis, we additionally report layer-specific weighted scores:
$\mathcal{P}=\frac{\sum_{i\in\mathcal{P}}w_{i}\cdot s_{i}}{\sum_{i\in\mathcal{P}}w_{i}}\quad\mathcal{M}=\frac{\sum_{i\in\mathcal{M}}w_{i}\cdot s_{i}}{\sum_{i\in\mathcal{M}}w_{i}}\quad\mathcal{R}=\frac{\sum_{i\in\mathcal{R}}w_{i}\cdot s_{i}}{\sum_{i\in\mathcal{R}}w_{i}}$ (7)
where $\mathcal{P}$, $\mathcal{M}$, and $\mathcal{R}$ refer to the sets of subtasks associated with the Perception, Memory, and Reasoning layers, respectively. This fine-grained evaluation enables targeted diagnosis and provides actionable insights for both model development and benchmark refinement.
A.6 In-depth Interpretation of Descriptive Statistics and Normality Tests
The descriptive statistics and normality test results for GoldBench latent variables reveal several important patterns in the MLLM’s score data distribution:
General Characteristics
The standardized latent scores maintain a mean of approximately zero and unit variance, confirming proper standardization across all dimensions. However, the observed score ranges exhibit moderate variation (-3.68 to 2.37), with Perception showing the widest spread of values.
Distributional Properties
  • Skewness Patterns: The Perception dimension shows significant left-skewness ($\gamma_{1}=-0.974$), indicating a longer left tail, while Reasoning exhibits slight right-skewness ($\gamma_{1}=0.399$). Memory and GoldBench overall show negligible skewness ($|\gamma_{1}|<0.5$).
  • Kurtosis Characteristics: Perception demonstrates marked leptokurtosis ($\gamma_{2}=1.075$), suggesting a more peaked distribution with heavier tails compared to the normal distribution. All other dimensions display mesokurtic or slightly platykurtic tendencies.
Normality Assessment
The Cramér-von Mises test reveals critical differences in distributional conformity:
  • Memory maintains normality ($p=0.388$) with minimal deviation from theoretical expectations.
  • GoldBench overall ($p=0.007$), Perception ($p=0.001$), and Reasoning ($p=0.006$) all significantly depart from normality, particularly through combined skewness and kurtosis violations.
Table 4: Descriptive statistics and Cramér-von Mises normality test for GoldBench latent variables.
Latent Variable Mean Median Min Max Std. Dev Excess Kurtosis Skewness N C-vM Statistic p-value
GoldBench -0.000 0.032 -2.687 1.954 1.000 0.021 -0.510 190 0.192 0.007
Memory -0.000 0.093 -2.994 2.341 1.000 0.066 -0.333 190 0.059 0.388
Perception 0.000 0.121 -3.684 1.484 1.000 1.075 -0.974 190 0.270 0.001
Reasoning -$0$0.100 -1.739 2.370 1.000 -0.709 0.399 190 0.193 0.006
A.7 SEM-based Evaluation Analysis of MMEBench
We apply our SEM-based capability modeling framework to analyze the internal structure of the MMEBench benchmark. Using performance data from 191 MLLMs, we construct a hierarchical model where individual tasks serve as observed indicators grouped into latent cognitive dimensions such as Coarse-Grained Tasks (C_G), Fine-Grained Tasks (F_G) and Reasoning Tasks capabilities. These first-order latent variables collectively form a higher-order cognition construct.
Table 5 summarizes the descriptive statistics and normality test results of the latent variables extracted from MMEBench.
Table 5: Descriptive statistics and Cramér-von Mises normality test for MMBench latent variables.
Latent Variable Mean Median Min Max Std. Dev Excess Kurtosis Skewness N C-vM Statistic p-value
C_G 0.000 0.220 -4.634 1.096 1.000 6.066 -2.208 191 1.647 0.000
F_G 0.000 0.190 -4.589 1.407 1.000 3.921 -1.645 191 0.770 0.000
MME 0.000 $0$3.006 2.194 1.000 0.385 -0.506 191 0.145 0.027
OCR -0.000 0.205 -3.160 1.701 1.000 -0.626 -0.416 191 0.417 0.000
cognition $0$0.180 -3.021 1.991 1.000 -0.492 0.107 191 0.410 0.000
The latent variables were standardized during analysis, yielding mean values near zero and unit standard deviations across all constructs. Observed score ranges showed considerable spread, with minimum values spanning from -4.63 to -3.00 and maximum values between 1.10 to 2.19, reflecting natural variation in the underlying traits.
Distributional characteristics revealed non-normal patterns: Skewness coefficients ranged from -2.21 (moderate left skew) to 0.11 (near-symmetry), while excess kurtosis values varied from -0.63 (mild platykurtosis) to 6.07 (pronounced leptokurtosis). Notably, construct C_G demonstrated particularly heavy tails combined with negative skewness, suggesting atypical response patterns in this domain.
Formal normality testing using the Cramér-von Mises criterion produced statistically significant results (all p < 0.05), rejecting the null hypothesis of normal distributions. This non-normality persisted despite standardization, likely reflecting true population characteristics rather than measurement artifacts. The observed distributional features - particularly the combination of negative skewness and elevated kurtosis in some constructs - warrant consideration when interpreting structural equation modeling outputs, as standard maximum likelihood estimation assumes multivariate normality.
Table 6: Outer loadings of MMEBench tasks on their respective latent cognitive dimensions. Values above 0.70 indicate strong contributions.
Indicator Latent Variable Outer Loading
Artwork F_G 0.826
Celebrity F_G 0.768
Code Reasoning cognition 0.930
Color C_G 0.939
Commonsense Reasoning cognition 0.914
Count C_G 0.938
Existence C_G 0.708
Landmark F_G 0.811
Numerical Calculation cognition 0.939
OCR OCR 1.000
Position C_G 0.851
Posters F_G 0.933
Scene F_G 0.618
Text Translation cognition 0.743
The outer loadings (Table 6) indicate strong task contributions to their respective cognitive dimensions, with most loadings exceeding 0.70 and several above 0.90. This demonstrates good indicator reliability and that the selected tasks effectively measure their intended latent abilities. However, there are still some tasks that do not contribute much, which gives us room for further optimization.
Table 7: Variance Inflation Factor (VIF) values for MMEBench tasks indicating multicollinearity among observed indicators.
Task VIF
Artwork 2.590
Celebrity 1.819
Code Reasoning 4.967
Color 4.262
Commonsense Reasoning 2.824
Count 3.432
Existence 1.880
Landmark 2.900
Numerical Calculation 4.714
OCR 1.000
Position 2.341
Posters 2.128
Scene 1.698
Text Translation 2.105
However, the Variance Inflation Factor (VIF) values reported in Table 7 reveal moderate multicollinearity among tasks. While most VIF values remain below the common threshold of 5, certain tasks such as Code Reasoning (4.967) and Numerical Calculation (4.714) approach this limit, suggesting some redundancy that could potentially undermine measurement efficiency.
Table 8: HTMT matrix showing discriminant validity among MMEBench latent constructs. Values above 0.90 indicate potential overlap.
C_G F_G MME OCR cognition
C_G - - - - -
F_G 0.910 - - - -
MME 0.357 0.317 - - -
OCR 0.491 0.444 0.499 - -
cognition 0.725 0.690 0.542 0.627 -
Furthermore, the Heterotrait-Monotrait (HTMT) ratio matrix in Table 8 exposes limited discriminant validity between latent constructs. Several HTMT values, particularly between Coarse-Grained Tasks (C_G) and Fine-Grained Tasks (F_G) at 0.91, exceed the recommended cutoff of 0.90. This indicates conceptual overlap between these ability dimensions and highlights a need for clearer distinction in benchmark task design.
Overall, while MMEBench tasks reliably measure their target cognitive dimensions, the presence of task redundancy and insufficient discriminant validity suggest that future benchmark iterations should refine task selection to reduce overlap and enhance construct clarity.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main contributions of this paper are presented in a list at the end of the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We state potential limitations in the appendices.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We describe the algorithm and what the different symbols refer to in order to ensure that reviewers understand the implementation details.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide sample data in the article and will open source all code and data.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide instructions on data access and preparation in data curation process and will open source all data.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The full details of our experimental setting is demonstrated in Section 5.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have presented the relevant statistical analysis in the form of graphs and charts in the main text.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We state the detailed size of the proposed dataset and the time cost of inference is related to the computational resources used.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All images containing sensitive personainformation, such as visible faces, identifiable personal details, or private life scenarios, wereeither excluded or processed to blur or mask sensitive regions, to mitigate privacy risks.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our data are collected from existing dataset and desensitized, no personal privacy or social security is involved.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We describe a detailed data collection process and privacy protection strategy, with all data coming from existing dataset.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original paper that produced the code package or dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide a new recollection dataset with CC-BY 4.0 license.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our core method is based on a multimodal large language model.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.